# Real-time Tracking in a Status Update System with an Imperfect Feedback Channel

### v1

@@ III-B Belief-MDP Problem
Due to the imperfect feedback channel, the estimated value of the source state at the sink $\hat{X}(t)$ is only partially observable at the transmitter side. Consequently, only the observable part of the state $o(t)$ is accessible for decision-making, which is incomplete state information and poses challenges for (optimal) transmission scheduling. To address the information insufficiency, we utilize the notion of belief [33, Chapter 7] and cast the POMDP defined in Section III-A as a belief-MDP. To this end, we define state-like quantities, referred to as belief-states, that retain the Markov property and capture all information necessary for optimal transmission scheduling. The belief-state consists of the observable component $o(t)$ of the state and a belief about the partially observable component $\hat{X}(t)$.
Let $I^{c}(t)$ denote the complete information state at slot $t$, which consists of: 1) initial probability distribution over the states, 2) the history of past and current observations, i.e., ${\{o(1),\dots,o(t)\}}$, and 3) the history of past actions, i.e., ${\{a(1),\dots,a(t-1)\}}$. Let $\rho(t)$ denote the belief about $\hat{X}(t)$ at slot $t$. The belief is a $N$-dimensional vector ${\rho(t)=[\rho_{1}(t),\dots,\rho_{N}]}$ that the entries are the conditional probability of the possible values of $\hat{X}(t)$ given $I^{c}(t)$. Formally, the $i$th entry of the belief is given as ${\rho_{i}(t)=\mathrm{Pr}\big{(}\hat{X}(t)=i\mid I^{c}(t)\big{)}}$.
@@
  • Case2: $a(t)=1$ and $d(t)=1$. The transmitter sends a status update containing the information $X(t)$ and receives ACK. In this case, the transmitter infers the exact value of sink’s estimate at slot $t+1$, i.e., $\hat{X}(t+1)=X(t)$. Thus, the entry of $\rho$ associated with $X(t)$ becomes one, and all others become zero. The belief update is given as
    $\rho_{i}(t+1)=\begin{cases}1&i=X(t)\\ 0&i\neq X(t),\end{cases}$ (9)
    which can be written compactly as $\rho(t+1)=e_{X(t)}$ where $e_{n}$ is the $n$th column of identity matrix $I_{N}$. We refer to this belief as reset belief to $X(t)$.
  • Case3: $a(t)=1$ and $d(t)=0$. The transmitter sends the status update $X(t)$ but does not receive ACK. This can be caused by one of the following two events: 1) event $E_{1}$ where the status update is received at the sink but ACK is lost over the feedback channel, and 2) event $E_{2}$ where the sink does not receive the status update due to an error over the forward channel. Under $E_{1}$, sink’s estimate at slot $t+1$ become $\hat{X}(t+1)=X(t)$; under $E_{2}$, sink’s estimate remains unchanged, i.e., $\hat{X}(t+1)=\hat{X}(t)$. Thus, the belief update can be written as
@@
    $\rho_{i}(t+1)=\begin{cases}\displaystyle\frac{p_{s}(1-p_{f})}{1-p_{s}p_{f}}+\frac{1-p_{s}}{1-p_{s}p_{f}}\rho_{i}(t)&i=X(t)\\ \displaystyle\frac{1-p_{s}}{1-p_{s}p_{f}}\rho_{i}(t)&i\neq X(t).\end{cases}$ (13)
@@
Belif-states: Let ${l(t)\!=\!\big{(}X(t),b(t),\rho(t),d(t\!-\!1),X(t\!-\!1)\big{)}\!\in\!\mathcal{L}}$ denote the belief-state at slot $t$, where $\mathcal{L}$ is the belief-state space.
Actions: Let $a(t)\in\mathcal{A}_{l}$ denote the action at slot $t$, where $\mathcal{A}_{l}$ is the space of the feasible actions at belief-state $l$, defined as
$\mathcal{A}_{l}=\big{\{}a\in\{0,1\}\mid b-a\geq 0\big{\}}.$ (14)
Note that since the action space depends on the observable part of the belief-state $o(t)$, the action space in the POMDP and the belief-MDP are the same.
@@
$C\big{(}l(t)\big{)}=\sum_{i=1}^{N}\rho_{i}(t)g\big{(}X(t),i\big{)}.$ (15)
Belief-state transition probabilities: Let ${\mathrm{Pr}(l^{\prime}\mid l,a)}$ denote the probability of moving from the current belief-sate ${l=\big{(}X,b,\rho,d_{-1},X_{-1}\big{)}}$ to the next belief-state ${l^{\prime}=\big{(}X^{\prime},b^{\prime},\rho^{\prime},d_{-1}^{\prime},X_{-1}^{\prime}\big{)}}$ under action $a$. For the sake of brevity, we use $\tilde{\rho}$ to denote the value of the next belief for the case $a(t)=1$ and $d(t)=0$, obtained from (13). The belief-state transition probabilities can be expressed as
$\mathrm{Pr}(l^{\prime}\mid l,a)=$
@@
Note that, for a general MDP, the optimal value $C^{*}$ and the optimal policies $\pi^{*}$ may depend on the initial state $l(0)$. However, Theorem 1 below shows that the belief-MDP (17) is communicating [34, Chapter 8.3.1] and, therefore, its optimal value $C^{*}$ is independent of the initial state.
@@
Follows directly from [35, Proposition 4.2.3]. ∎
@@
$\bar{h}+h(l)=\min_{a\in\mathcal{A}_{l}}\left[C(l)+\sum_{l^{\prime}\in\mathcal{L}}\mathrm{Pr}(l^{\prime}\mid l,a)h(l^{\prime})\right],\leavevmode\nobreak\ \forall l\in\mathcal{L},$ (18)
@@
Follows directly from [35, Proposition 4.2.3]. ∎
@@
III-C RVIA-based Solution to Belief-MDP Problem
The relative value iteration algorithm (RVIA) [35, Section 4.3] turns Bellman’s equation (18) into an iterative procedure which selects an arbitrary reference state $l_{\text{ref}}\in\mathcal{L}$, kept fixed throughout the iterations, and each step $i$ updates the estimates $\{h^{i}(l)\}_{l\in\mathcal{L},l\neq l_{\text{ref}}}$ of the relative value function, until convergence. Specifically, we initialize $h^{0}(l)=0$ for all $l\in\mathcal{L}$ and for $i=1,2,\ldots$, we update
$V^{i}(l)$ $=\min_{a\in\mathcal{A}_{l}}\left[C(l)+\sum_{l^{\prime}\in\mathcal{L}}\mathrm{\Pr}(l^{\prime}\mid l,a)h^{i-1}(l^{\prime})\right],$ (19)
@@ III-C RVIA-based Solution to Belief-MDP Problem
To employ RVIA, the belief-state space (and the action space) must be finite. However, since the believe entry $\rho(t)$ is continuous, i.e., $\rho(t)\in\mathbb{R}$, the belief-state space $\cal{L}$ become infinite. Fortunately, the belief-evolution follows a specific pattern which we exploit to truncate the belief-state space $\cal{L}$ and subsequently to develop a practical approximate solution using RVIA.
To express a belief space compactly let $u_{1}$ and $u_{2}$ denote the probabilities in (10) as ${u_{1}=\mathrm{Pr}(E_{1}\mid\overline{\text{ACK}})}$ and ${u_{2}=\mathrm{Pr}(E_{2}\mid\overline{\text{ACK}})}$. Note that, whenever the transmitter sends a status update and receives ACK, the belief is set to a specific value given by (9) (i.e., $e_{X(t)}$). Furthermore, the belief remains unchanged whenever the transmitter stays idle (see (8)). Thus, from (10), the space of all possible values that $\rho(t)$ can reach during $m$ consecutive transmission attempts without receiving ACK can be expressed as
$\Omega_{m}=\cup_{i=1}^{m}\Phi(i),$ (21)
where
@@
To obtain a finite belief-state space $\cal L$, we restrict all the possible values of the belief $\rho(t)$ to those contained in ${\Omega}_{m}$. This is reasonable because the probability of not receiving feedback during $m$ consecutive transmission attempts is $(1-p_{s}p_{f})^{m}$, i.e., it decreases exponentially in $m$. Thus, for a large enough $m$, the space of all possible values of $\rho(t)$ is well approximated by ${\Omega}_{m}$. In the unlikely (but still possible) event when the number of consecutive unsuccessful transmission attempts exceeds $m$, we replace the value of the belief $\rho(t)$ obtained from (10) by the closest value contained in set ${\Omega}_{m}$. Closeness is measured using the Kullback–Leibler (KL) divergence [36, Chapter 2]. We define the updated belief $\rho^{*}$ as
@@
where the KL divergence between the beliefs is given by
$\operatorname{KL}\left(\rho(t)||\rho\right)=\sum_{i=1}^{N}\rho_{i}(t)\log\frac{\rho_{i}(t)}{\rho_{i}}.$ (24)
This operation can be implemented efficiently via an offline generated look-up-table which provides the closest belief inside ${\Omega}_{m}$ for each element in $\Phi(m+1)$.
Replacing the beliefs of the belief-states associated with $\Phi(m+1)$ with the closest beliefs inside $\Omega_{m}$ adds new (possible) transitions between belief-states in the truncated belief-state space. Thus, we shall update the belief-state transition probabilities, i.e., $\Pr\{l^{\prime}|l,a\}$ given in (III-B), with a new state transition probabilities denoted as $\widetilde{\Pr}\{l^{\prime}|l,a\}$ for all $l^{\prime}$ and $l$ in the truncated belief-state space. A transition from $l\in\mathcal{L}$ to $l^{\prime}\in\mathcal{L}$ could possibly be directly, or indirectly by the projection, i.e., $l$ transits to a belief-state outside of the truncated belief-state space whose projection is $l^{\prime}$. Thus,
$\displaystyle\widetilde{\Pr}\{l^{\prime}\,|\,l,a\}={\Pr}\{l^{\prime}\,|\,l,a\}+\sum_{\underline{l}\in\underline{\mathcal{L}}}{\Pr}\{\underline{l}\,|\,l,a\},$ (25)
@@
Using truncated belief space and the transition probabilities in (25), we implement RVIA to solve the belief-MDP problem. The details of the proposed RVIA-based transmission policy are presented in Algorithm 1.
@@ 2       $i=i+1$
4             $V^{i}(l)=\min_{a\in\mathcal{A}_{l}}\left[C(l)+\sum_{l^{\prime}\in\mathcal{L}}\widetilde{\Pr}(l^{\prime}\mid l,a)h^{i-1}(l^{\prime})\right]$
@@ 5             $h^{i}(l)=V^{i}(l)-V^{i}(l_{\mathrm{ref}})$
Output: RVIA-based transmission policy $\pi^{*}=\pi$, the optimal value $C^{*}=V(l_{\mathrm{ref}})$
Algorithm 1 The RVIA-based transmission policy
III-D Low-complexity Solution to Belief-MDP Problem
The RVIA-based policy, presented in Algorithm 1, needs to explore all states and actions, and the size of the belief-state space grows exponentially with the number of source states and the battery’s capacity. Therefore, we develop a low-complexity dynamic transmission policy in which the belief-MDP problem is transformed into a sequence of per-slot optimization problems. In particular, at each slot $t$, the aim is to solve the following problem
@@ III-D Low-complexity Solution to Belief-MDP Problem
where ${\mathbb{E}\{C(t+1)\mid l(t)\}}$ is the (conditional) expected cost at slot $t+1$ given by (III-D), whose derivation is presented in Appendix D.
$\mathbb{E}\{C(t+1)\mid l(t)\}=g\big{(}X(t),X(t)\big{)}pp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)p(1-p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{j=1,j\neq X(t)}^{N}g\big{(}j,X(t)\big{)}(1-p)p_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)(1\!-\!p)(1\!-\!p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}g\big{(}X(t),i\big{)}\rho_{i}(t)(1-p)(1-\mathbb{E}\{a(t)\mid l(t)\})$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}g(j,i)\rho_{i}(t)p(1-\mathbb{E}\{a(t)\mid l(t)\}).$ (27)
We minimize (III-D) following the approach of opportunistically minimizing a (conditional) expectation [37], i.e., (III-D) is minimized by ignoring the expectations in each slot. To this end, we used the exhaustive search method. Note that since at each slot there are two possible actions, the exhaustive search method is an appropriate method to solve the per-slot optimization problem. Note that this policy does not require the truncation of the belief-state space.

### v2

@@ III-B Belief-MDP Problem
Due to the imperfect feedback channel, the estimated value of the source state at the sink $\hat{X}(t)$ is only partially observable at the transmitter side. Consequently, only the observable part of the state $o(t)$ is accessible for decision-making, which is incomplete state information and poses challenges for (optimal) transmission scheduling. To address the information insufficiency, we utilize the notion of belief [37, Chapter 7] and cast the POMDP defined in Section III-A as a belief-MDP. To this end, we define state-like quantities, referred to as belief-states, that retain the Markov property and capture all information necessary for optimal transmission scheduling. The belief-state consists of the observable component $o(t)$ of the state and a belief about the partially observable component $\hat{X}(t)$.
Let $I^{c}(t)$ denote the complete information state at slot $t$, which consists of: 1) initial probability distribution over the states, 2) the history of past and current observations, i.e., ${\{o(1),\dots,o(t)\}}$, and 3) the history of past actions, i.e., ${\{a(1),\dots,a(t-1)\}}$. Let $\rho(t)$ denote the belief about $\hat{X}(t)$ at slot $t$. The belief is a $N$-dimensional vector ${\rho(t)=[\rho_{1}(t),\dots,\rho_{N}(t)]}$ that the entries are the conditional probability of the possible values of $\hat{X}(t)$ given $I^{c}(t)$. Formally, the $i$th entry of the belief is given as ${\rho_{i}(t)=\mathrm{Pr}\big{(}\hat{X}(t)=i\mid I^{c}(t)\big{)}}$.
@@
  • Case 2: $a(t)=1$ and $f(t)=1$. The transmitter sends a status update containing the information $X(t)$ and receives ACK. In this case, the transmitter infers the exact value of sink’s estimate at slot $t+1$, i.e., $\hat{X}(t+1)=X(t)$. Thus, the entry of $\rho$ associated with $X(t)$ becomes one, and all others become zero. The belief update is given as
    $\rho(t+1)=e_{X(t)},$ (7)
    where $e_{n}$ is the $n$th column of identity matrix $I_{N}$. We refer to this belief as reset belief to $X(t)$.
  • Case 3: $a(t)=1$ and $f(t)=0$. The transmitter sends the status update $X(t)$ but does not receive ACK. This can be caused by one of the following two events: 1) event $E_{1}$ where the status update is received at the sink but ACK is lost over the feedback channel, and 2) event $E_{2}$ where the sink does not receive the status update due to an error over the forward channel. Under $E_{1}$, sink’s estimate at slot $t+1$ become $\hat{X}(t+1)=X(t)$; under $E_{2}$, sink’s estimate remains unchanged, i.e., $\hat{X}(t+1)=\hat{X}(t)$. Thus, the belief update can be written as
@@
    $\rho(t+1)=\frac{p_{s}(1-p_{f})}{1-p_{s}p_{f}}e_{X(t)}+\frac{1-p_{s}}{1-p_{s}p_{f}}\rho(t).$ (11)
@@
Belief-states: Let $l(t)\!=\!\big{(}X(t),b(t),\rho(t),f(t\!-\!1),X(t\!-\!1)\big{)}\!\in\!\mathcal{L}$ denote the belief-state at slot $t$, where $\mathcal{L}$ is the belief-state space.
Actions: Let $a(t)\in\mathcal{A}_{l}$ denote the action at slot $t$, where $\mathcal{A}_{l}$ is the space of the feasible actions at belief-state $l$. Note that since the action space depends on the observable part of the belief-state $o(t)$, the action space in the POMDP and the belief-MDP are the same.
@@
$C\big{(}l(t)\big{)}=\sum_{i=1}^{N}\rho_{i}(t)d\big{(}X(t),i\big{)}.$ (12)
Belief-state transition probabilities: Let ${\mathcal{P}(l^{\prime}\mid l,a)}$ denote the state transition probability defined as the probability of moving from the current belief-sate ${l=\big{(}X,b,\rho,f_{-1},X_{-1}\big{)}}$ to the next belief-state ${l^{\prime}=\big{(}X^{\prime},b^{\prime},\rho^{\prime},f_{-1}^{\prime},X_{-1}^{\prime}\big{)}}$ under action $a$. For the sake of brevity, we use $\tilde{\rho}$ to denote the value of the next belief for the case $a(t)=1$ and $f(t)=0$, obtained from (11). The belief-state transition probabilities can be expressed as
$\mathcal{P}(l^{\prime}\mid l,a)=$
@@
Note that, for a general MDP, the optimal value $C^{*}$ and the optimal policies $\pi^{*}$ may depend on the initial state $l(0)$. However, Theorem 1 below shows that the belief-MDP (14) is communicating [38, Chapter 8.3.1] and, therefore, its optimal value $C^{*}$ is independent of the initial state.
@@
Follows directly from [39, Proposition 4.2.3]. ∎
@@
$\bar{h}\!+\!h(l)\!=\!\min_{a\in\mathcal{A}_{l}}\left[C(l)+\sum_{l^{\prime}\in\mathcal{L}}\mathcal{P}(l^{\prime}\mid l,a)h(l^{\prime})\right]\!,\leavevmode\nobreak\ \forall l\in\mathcal{L},$ (15)
@@
Follows directly from [39, Proposition 4.2.3]. ∎
@@
III-C POMDP-based Solution to Belief-MDP Problem
The relative value iteration algorithm (RVIA) [39, Section 4.3] turns Bellman’s equation (15) into an iterative procedure which selects an arbitrary reference state $l_{\text{ref}}\in\mathcal{L}$, kept fixed throughout the iterations, and each step $i$ updates the estimates $\{h^{i}(l)\}_{l\in\mathcal{L},l\neq l_{\text{ref}}}$ of the relative value function, until convergence. Specifically, we initialize $h^{0}(l)=0$ for all $l\in\mathcal{L}$ and for $i=1,2,\ldots$, we update
$V^{i}(l)$ $=\min_{a\in\mathcal{A}_{l}}\left[C(l)+\sum_{l^{\prime}\in\mathcal{L}}\mathcal{P}(l^{\prime}\mid l,a)h^{i-1}(l^{\prime})\right],$ (16)
@@ III-C POMDP-based Solution to Belief-MDP Problem
To employ RVIA, the belief-state space (and the action space) must be finite. However, since the believe entry $\rho(t)$ is continuous, i.e., $\rho(t)\in\mathbb{R}$, the belief-state space $\cal{L}$ become infinite. Fortunately, the belief evolution follows a specific pattern, which we exploit to truncate the belief-state space $\cal{L}$ and subsequently to develop a practical approximate solution using RVIA.
To express a belief space compactly let $u_{1}$ and $u_{2}$ denote the probabilities in (8) as ${u_{1}=\mathrm{Pr}(E_{1}\mid\overline{\text{ACK}})}$ and ${u_{2}=\mathrm{Pr}(E_{2}\mid\overline{\text{ACK}})}$. Note that, whenever the transmitter sends a status update and receives ACK, the belief is set to a specific value given by (7) (i.e., $e_{X(t)}$). Furthermore, the belief remains unchanged whenever the transmitter stays idle (see (6)). Thus, from (8), the space of all possible values that $\rho(t)$ can reach during $m$ consecutive transmission attempts without receiving ACK can be expressed as $\Omega_{m}=\cup_{i=1}^{m}\Phi(i)$, where
@@
To obtain a finite belief-state space $\cal L$, we restrict all the possible values of the belief $\rho(t)$ to those contained in ${\Omega}_{m}$. This is reasonable because the probability of not receiving feedback during $m$ consecutive transmission attempts is $(1-p_{s}p_{f})^{m}$, i.e., it decreases exponentially in $m$. Thus, for a large enough $m$, the space of all possible values of $\rho(t)$ is well approximated by ${\Omega}_{m}$. In the unlikely (but still possible) event when the number of consecutive unsuccessful transmission attempts exceeds $m$, we replace the value of the belief $\rho(t)$ obtained from (8) by the closest value contained in set ${\Omega}_{m}$. Closeness is measured using the Kullback–Leibler (KL) divergence [40, Chapter 2]. We define the updated belief $\rho^{*}$ as
@@
where the KL divergence between the beliefs is given by $\operatorname{KL}\left(\rho(t)||\rho\right)=\sum_{i=1}^{N}\rho_{i}(t)\log\frac{\rho_{i}(t)}{\rho_{i}}$. This operation can be implemented efficiently via an offline generated look-up-table which provides the closest belief inside ${\Omega}_{m}$ for each element in $\Phi(m+1)$.
Replacing the beliefs of the belief-states associated with $\Phi(m+1)$ with the closest beliefs inside $\Omega_{m}$ adds new (possible) transitions between belief-states in the truncated belief-state space. Thus, we shall update the belief-state transition probabilities, i.e., $\mathcal{P}(l^{\prime}\mid l,a)$ given in (III-B), with a new state transition probabilities denoted as $\widetilde{\mathcal{P}}(l^{\prime}\mid l,a)$ for all $l^{\prime}$ and $l$ in the truncated belief-state space. A transition from $l\in\mathcal{L}$ to $l^{\prime}\in\mathcal{L}$ could possibly be directly, or indirectly by the projection, i.e., $l$ transits to a belief-state outside of the truncated belief-state space whose projection is $l^{\prime}$. Thus,
$\displaystyle\widetilde{\mathcal{P}}(l^{\prime}\mid l,a)=\mathcal{P}(l^{\prime}\mid l,a)+\sum_{\underline{l}\in\underline{\mathcal{L}}}{\Pr}\{\underline{l}\,|\,l,a\},$ (20)
@@
Using truncated belief space and the transition probabilities in (20), we implement RVIA to solve the belief-MDP problem. The details of the proposed POMDP-based transmission policy are presented in Algorithm 1.
@@ 2       $i=i+1$
4             $V^{i}(l)=\min_{a\in\mathcal{A}_{l}}\left[C(l)+\sum_{l^{\prime}\in\mathcal{L}}\widetilde{\mathcal{P}}(l^{\prime}\mid l,a)h^{i-1}(l^{\prime})\right]$
@@ 5             $h^{i}(l)=V^{i}(l)-V^{i}(l_{\mathrm{ref}})$
Output: POMDP-based transmission policy $\pi^{*}=\pi$, the optimal value $C^{*}=V(l_{\mathrm{ref}})$
Algorithm 1 The POMDP-based transmission policy
III-D Low-complexity Solution to the Belief-MDP Problem
The POMDP-based policy, presented in Algorithm 1, needs to explore all states and actions, and the size of the belief-state space grows exponentially with the number of source states and the battery’s capacity. Therefore, in the following, we introduce two low-complexity (LC) policies.
III-D1 Energy-Agnostic LC Policy
Inspired by the Lyapunov drift-plus penalty method [41], we develop an LC dynamic transmission policy in which the belief-MDP problem is transformed into a sequence of per-slot optimization problems. In particular, at each slot $t$, the aim is to solve the following problem
@@ III-D1 Energy-Agnostic LC Policy
where ${\mathbb{E}\{C(t+1)\mid l(t)\}}$ is the (conditional) expected cost at slot $t+1$ given by (III-D1), whose derivation is presented in Appendix D.
$\mathbb{E}\{C(t+1)\mid l(t)\}=d\big{(}X(t),X(t)\big{)}pp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)p(1-p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{j=1,j\neq X(t)}^{N}d\big{(}j,X(t)\big{)}qp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)q(1\!-\!p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)p(1-\mathbb{E}\{a(t)\mid l(t)\})$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)q(1-\mathbb{E}\{a(t)\mid l(t)\}).$ (22)
We minimize (III-D1) following the approach of opportunistically minimizing a (conditional) expectation [41], i.e., (III-D1) is minimized by dropping the expectations in each slot. Since there are only two possible actions at each slot, the exhaustive search method is an appropriate method to solve the per-slot optimization problem. If the resulting expected cost for both actions are the same, the action of staying idle, i.e., $a=0$, is selected. Note that this policy does not require the truncation of the belief-state space.
The above proposed policy is agnostic to the energy arrival process, while the energy arrival process influences the long-term availability of energy for the transmitter’s operation. Thus, ignoring the energy arrival rate, especially when the energy arrival rate is low, would lead to inefficient decision-making. To address this shortcoming, we propose an energy-aware LC policy, as detailed below.
III-D2 Energy-Aware LC Policy
To take the energy arrival process into account, we introduce a regularization term to add to the objective function of Problem (21). The regularization term is a function of the energy arrival process and is meant to promote less transmissions to save energy for future use. More precisely, the regularization term decreases the transmission rate when the energy arrival rate decreases, while its effect diminishes as the energy arrival rate increases. This term is defined as $-\gamma a(t)\big{(}e(t)-1\big{)}$, where $\gamma$ is a positive tuning parameter. Thus, the new per-slot optimization problem is formulated as
[…]
The conditional expectation ${\mathbb{E}\{-\gamma a(t)\big{(}e(t)-1\big{)}\mid l(t)\}}$ is given as
$\mathbb{E}\{-\gamma a(t)\big{(}e(t)-1\big{)}\mid l(t)\}=-\gamma\mathbb{E}\{a(t)\mid l(t)\}(\mu-1),$ (24)
where the equality followed from the fact that ${\mathbb{E}\{e(t)\mid l(t)\}=\mu}$, and the action decision at slot $t$ is independent of the energy arrival during the slot $t$. Thus, substituting (III-D1) and (24) into (23), the expected cost function is given as
$\mathbb{E}\{C(t+1)-\gamma a(t)\big{(}e(t)-1\big{)}\mid l(t)\}$
$=d\big{(}X(t),X(t)\big{)}pp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)p(1-p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{j=1,j\neq X(t)}^{N}d\big{(}j,X(t)\big{)}qp_{s}\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)q(1\!-\!p_{s})\mathbb{E}\{a(t)\mid l(t)\}$
$+\sum_{i=1}^{N}d\big{(}X(t),i\big{)}\rho_{i}(t)p(1-\mathbb{E}\{a(t)\mid l(t)\})$
$+\sum_{i=1}^{N}\sum_{j=1,j\neq X(t)}^{N}d(j,i)\rho_{i}(t)q(1-\mathbb{E}\{a(t)\mid l(t)\})$
$-\gamma\mathbb{E}\{a(t)\mid l(t)\}(\mu-1).$ (25)
We follow the same approach to minimize (24) as we employed to minimize (III-D1). In particular, the energy-agnostic LC policy is a special case of the energy-aware LC policy when the tuning parameter is set to zero, i.e., $\gamma=0$.